A total of 315 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold were two times the number of adult tickets sold. how many adult tickets were sold.

Answer by ikleyn(52915)

.
The condition says us that


Adult tickets + student tickets = 315,   or


A             + S               = 315.    (A stays for adults tickets, while S stays for the student ones).


It also says that


S = 2A.


So you can replace S in the very first equation by 2A. Then you get


A + 2A = 315,   or

3A = 315,


which implies  A = 315/3 = 105.


Answer.  105 adult tickets were sold.

Solved.